You are now asked to journalize and post the adjusting entries after completing the worksheet. This is because, in the manual accounting system, the worksheet is not part of the journals and ledgers. Therefore completing the Adjustments column in the worksheet does not automatically update the appropriate ledger accounts. Since this practice set is a manual accounting practice set, you are still required to manually journalize the adjusting entries and post them to the general ledger as a separate step.

The purpose of adjusting entries is to ensure that the general ledger account balances reflect the revenues earned and expenses incurred during the accounting period.

Remember that the details of the end of month adjustments for June are as follows:

  • Office Furniture owned by the business: original purchase price was $7,000, estimated useful life was 6 years, and estimated residual value was $1,000 at the end of the useful life. Depreciation is calculated on a monthly basis using the straight line method. The monthly depreciation charge is calculated as the yearly depreciation expense divided by the number of months in a year.
  • Office Equipment owned by the business: original purchase price was $51,000, estimated useful life was 10 years, and estimated residual value was $6,000 at the end of the useful life. Depreciation is calculated on a monthly basis using the straight line method. The monthly depreciation charge is calculated as the yearly depreciation expense divided by the number of months in a year.
  • The water usage for the month of June is estimated to be $1,155.
  • The estimated electricity payable as at the end of June is $303.
  • Sales staff work every single day during the week including weekends and are paid on a weekly basis. Wages were last paid up to and including 24 June. Wages incurred after that day (from June 25 to June 30 inclusive) are estimated to have been $310 per day.
  • Interest expense incurred during the month of June but not yet paid to Earth Bank for the bank loan is $200.
  • Office supplies totaling $3,944 are still on hand at June 30.
  • One month of prepaid rent expired during June. Remember that at the beginning of June, the business paid two month's worth of rent in advance.
  • 3 months of advertising remained pre-paid at the start of June.
  • 5 months of insurance remained pre-paid at the start of June.

When calculating the portion of prepayments that expire during the month of June, you are asked to assume that an equal amount of expense is incurred per month.

After taking a physical count of inventory at the end of the month, the balance of inventory on hand as at June 30 is found to be equal to the closing balance of the Merchandise Inventory account. This means there is no adjusting entry required for inventory shrinkage.

Instructions for adjusting entries

1)Use the worksheet to assist you in recording the adjusting entries in the general journal.

2)Post the adjusting entries from the general journal to the general ledger accounts.

3)Record the final adjusted balance of each ledger account in the Adjusted Balance row of each ledger, even for ledgers with a balance of zero. Although each ledger already has a running balance, the Adjusted Balance row must still be filled out in order to receive full points.

GENERAL JOURNAL Date Account and Explanation Post Ref. Debit Credit Jun 30 (select) (select) Jun 30 (select) (select) Jun 30GENERAL LEDGER Account: Office Supplies Account No. 130 Date Ref. Debit Credit Balance Description 31 Balance May 3,981 DR JuAccount: Accum Depn: Office Furniture Account No. 151 Date Description Ref. Debit Credit Balance May 31 Balance 1,083 CR JunAccount: Water Payable Account No. 222 Date Description Ref. Debit Credit Balance May 31 Balance 1,120 CR Jun 30 Jun 30 Jun 3Account: Wages Expense Account No. 516 Date Description Ref. Debit Credit Balance May 31 Balance 0 Jun 30 Total from cash payAccount: Insurance Expense Account No. 543 Date Description Ref. Debit Credit Balance May 31 Balance 0 Jun 30 Jun 30 oud JunAccount: Interest Expense Account No. 571 Date Ref. Debit Credit Balance Description 31 Balance May 0 Jun 30 4 Jun 30 Jun 30
Answer #1

1. Dep of office furn

Depreciation p.a. = ($7,000 - $1,000) / 6 = $1,000

Depreciation per month = $1,000/12 = $83

Depreciation: Office Furn Expense Dr. 83

To Accumulated Depn: Off furn 83

2. Dep of office equipment

Depreciation p.a. = ($51,000 - $6,000) / 10 = $4,500

Depreciation per month = $4,500/12 = $375

Depreciation: Office Equip Expense Dr. 375

To Accumulated Depn: Off Equip 375

3. Water Expense Dr. 1,155

To Water Payable 1,155

4. Electricity Expense Dr. 303

To Electricity Payable 303

5. Wages Expense Dr 1,860

To Wages Payable 1,860

$310 per day * 6 days = $1,860

6. Interest Expense Dr. 200

To Interest Payable 200

7. Office Supplies expense = $4,382 - $3944 = $438

Office Supplies expense Dr. 438

To Office Supplies 438

8. Rent Expense Dr. 4,750

To Prepaid Rent 4,750

Expense for June = 9500/2

9. Advertisement Expense Dr. 4,000

To Prepaid Advertisement 4,000

Expense for June = 12,000/3 = 4,000

10. Insurance Expense Dr. 840

To Prepaid Insurance 840

Expense for June = 4,200 /5 = 840

Office Supplies

Date Description Ref. Debit Credit Balance
Office Supplies Expense 438
Balance 3944

Prepaid Rent

Date Description Ref. Debit Credit Balance
Rent Expense 4750

Prepaid Advertising

Date Description Ref. Debit Credit Balance
Advertisement Expense 4000

Prepaid Insurance

Date Description Ref. Debit Credit Balance
Insurance Expense 840

Accumulated Depreciation : Office Furniture

Date Description Ref. Debit Credit Balance
Depreciation : Office Furniture Expense 83

Accumulated Depreciation : Office Equipment

Date Description Ref. Debit Credit Balance
Depreciation : Office Equipment Expense 375

Wages Payable

Date Description Ref. Debit Credit Balance
Wages Expense 1860

Electricity Payable

Date Description Ref. Debit Credit Balance
Electricity Expense 303

Water Payable

Date Description Ref. Debit Credit Balance
Water Expense 1155

Interest Payable

Date Description Ref. Debit Credit Balance
Interest Expense 200

Interest Revenue

Date Description Ref. Debit Credit Balance

Advertising Expense

Date Description Ref. Debit Credit Balance
Prepaid Advertising 4000

Wages Expense

Date Description Ref. Debit Credit Balance
Wages Payable 1860

Rent Expense

Date Description Ref. Debit Credit Balance
Prepaid Rent 4750

Electricity Expense

Date Description Ref. Debit Credit Balance
Electricity Payable 303

Water Expense

Date Description Ref. Debit Credit Balance
Water Payable 1155

Insurance Expense

Date Description Ref. Debit Credit Balance
Prepaid Insurance 840

Office Supplies Expense

Date Description Ref. Debit Credit Balance
Office Supplies 438

Depreciation : Office Furniture Expense

Date Description Ref. Debit Credit Balance
Accumulated Depreciation : Office Furniture 83

Depreciation : Office Equipment Expense

Date Description Ref. Debit Credit Balance
Accumulated Depreciation : Office Equipment 375

Interest Expense

Date Description Ref. Debit Credit Balance
Interest Payable 200

All the entries are of June 30 and all adjusting entries as required are posted to T accounts